OpenStudy (anonymous):

The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

@Mr.ClayLordMath Help?

OpenStudy (anonymous):

@sourwing Help?

OpenStudy (whpalmer4):

what's the area of a rectangle?

OpenStudy (whpalmer4):

Area of a rectangle is \[A = l*w\]where \(w\) is the width and \(l\) is the length. We know the area of this rectangle is 280, so \(280 = l * w\). Next, we know that if we change the length to be 5 meters less, and the width to be 1 meter more, they will be equal (because the shape will be a square). That implies that \[l-5 = w+1\] Using the two equations \[280 = l * w\]\[l-5=w+1\]you can work out the values of the length and width.

OpenStudy (whpalmer4):

I would suggest an easy approach might be to solve the second equation for \(l\), then plug whatever that expression is in place of \(l\) in the first equation. That will give you a formula in terms of \(w\) which you can solve. Note that you'll get two solutions, as it is a quadratic equation, but only one will make sense in this problem.